How would I determinethe maximum number of letres of milk (with a density of 1030Kg m^-3 ) that can be added to a 50L drum that weights 70N when empty, so that the drum will float in fresh water (density 1000 Kg m^-3)

Grade:

2 Answers

bharat bajaj IIT Delhi
9 years ago
If the drum would float in water, then its density should be less than that of water
p < 1000
Say x litres of milk can be added.
Mass of drum = 70/10+ 1.03x ( Here x is in litres. Convert units)
Volume of drum = 50
p = M/V = (7 + 1.03x)/(50/1000) < 1000
7 + 1.03x < 50
x < 43/1.03
This is the maximum milk which you can add.
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty
KAPIL MANDAL
9 years ago
Thank you sir for this answer. Thank you very much.